OpenStudy (anonymous):

Find the equation for the line passing through (−1,−4) and 1,−17).

OpenStudy (anonymous):

the equation of line is as:- y-(-4)=(slope)(x-(-1)) y+4=(slope)(x+1) where slope = (-17-(-4))/(1-(-1)) = -13/2 y+4= (-13/2)(x+1) 2y+8=-13x-13 13x+2y+21=0